### v1

# Where Do Large Language Models Fail When Generating Code?

Abstract.
Large Language Models (LLMs) have shown great potential in code generation. However, current LLMs still cannot reliably generate correct code. Moreover, it is unclear what kinds of code generation errors LLMs can make. To address this, we conducted an empirical study to analyze incorrect code snippets generated by six popular LLMs on the HumanEval dataset. We analyzed these errors alongside two dimensions of error characteristics—semantic characteristics and syntactic characteristics—to derive a comprehensive code generation error taxonomy for LLMs through open coding and thematic analysis. We then labeled all 558 incorrect code snippets based on this taxonomy. Our results showed that the six LLMs exhibited different distributions of semantic and syntactic characteristics. Furthermore, we analyzed the correlation between different error characteristics and factors such as prompt length, code length, and test-pass rate. Finally, we highlight the challenges that LLMs may encounter when generating code and propose implications for future research on reliable code generation with LLMs.
1. Introduction
Automatically generating code from natural language has been a long-term pursuit across multiple research communities. Recent advances in Large Language Models (LLMs) have led to rapid, unprecedented improvements on this task (Nijkamp et al., 2022; Fried et al., 2022; Li et al., 2023a; OpenAI, 2023; cha, 2023). Despite this great progress, LLMs still cannot reliably generate correct code for many tasks. Currently, there is a lack of deep understanding of the cases where LLMs fail. Specifically, it remains unclear what types of code generation errors an LLM typically produces and whether different LLMs make similar errors. Answering these questions would help researchers gain insights into the limitations of existing models and identify opportunities for model improvement.
To bridge this knowledge gap, we conducted an in-depth analysis of code generation errors made by LLMs. We focused on six popular LLMs: CodeGen-16B (Nijkamp et al., 2022), InCoder-1.3B (Fried et al., 2022), GPT-3.5 (cha, 2023), GPT-4 (OpenAI, 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023a). These models produced 558 incorrect code solutions on the 164 tasks from the HumanEval dataset (Chen et al., 2021). Four authors performed multiple rounds of open coding and iterative refinement on these errors to derive a taxonomy of code generation errors made by LLMs. Specifically, we analyzed these errors alongside two dimensions—the semantic characteristics and syntactic characteristics of these errors:
Semantic characteristics can help identify the high-level root causes of these code generation errors. Representative semantic characteristics include missing condition, wrong (logical) direction, incorrect condition, etc. Analyzing these semantic characteristics across various LLMs can help understand the limitations of current LLMs in interpreting task requirements and generating semantically correct programs.
Syntactic characteristics can help localize where the error occurs in an incorrect code solution. Representative syntactic characteristics include incorrect code blocks, incorrect function arguments, incorrect return values, etc. Understanding these characteristics allows for a better assessment of current LLMs’ abilities to generate different kinds of code constructs. It can also help inform the design of new techniques for localizing and repairing code generation errors made by LLMs.
Our analysis shows that these LLMs tend to make errors with more complex syntactic characteristics, such as missing code blocks and incorrect code blocks, than smaller issues, such as incorrect method call target. We also found that the semantic characteristics behind these errors are likely to be intricate issues such as wrong (logical) direction and missing multiple steps, which implies a common challenge of interpreting relatively complex task descriptions for LLMs. These errors often require substantial code restructuring and repair rather than simple fixes. Furthermore, we observed different distributions of semantic and syntactic characteristics of code generation errors among the six LLMs, revealing the challenges of handling the diverse set of errors made by different LLMs.
In summary, this paper makes the following contributions:
We established a taxonomy of both syntactic and semantic characteristics of code generation errors through open coding and thematic analysis.
We analyzed the similarities and differences in errors made by different code generation models, highlighting the challenges faced by LLMs.
We discussed the implications and future opportunities for improving LLMs for code generation.
We developed an interactive data analysis website to help researchers and developers examine and explore code generation errors in different categories. The website is available at https://llm-code-errors.cs.purdue.edu.
2. Methodology
In this section, we first introduce the research questions of this study, followed by the descriptions of the dataset and LLMs used in our study and the procedure of our manual analysis.
2.1. Research Questions
This study investigates the following research questions.
RQ1: What specific types of errors manifest in code generation across different LLMs? This question aims to uncover the common characteristics and distinctions of different types of code generation errors. This can help us identify whether different LLMs exhibit unique or similar characteristics in their code generation errors.
RQ2: How much repair effort is needed for the incorrect code generated by LLMs? This question aims to measure how closely LLMs can mimic correct coding solutions and how many edits are needed to fix the errors. It can help us understand the potential difficulties of locating and repairing the incorrect model-generated code.
RQ3: How does the prompt complexity affect an LLM’s code generation? Prompting is an essential step in code generation with LLMs. This question aims to investigate how prompt complexity affects the correctness of code snippets generated by LLMs. It specifically examines how variations in prompt length might influence the types of errors in the generated code.
RQ4: How does code length relate to the types of LLMs’ code generation errors? Different from RQ5, this question aims to investigate if the generated code length affects the correctness of a code snippet generated by LLMs. We are particularly interested in those lengthy code solutions and their error characteristics.
RQ5: Does partially failed code exhibit different characteristics compared with fully failed code? This question explores the distinctions between code that fails a subset of test cases and code that fails all test cases. It can offer insights into the specific challenges faced in achieving full correctness on test cases.
2.2. Code Generation LLMs
In this study, we focus on six representative code generation LLMs: CodeGen-16B (Nijkamp et al., 2022), InCoder-1.3B (Fried et al., 2022), GPT-3.5 (cha, 2023), GPT-4 (OpenAI, 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023a). As shown in Table 1, these models cover a wide range of model sizes (1B to 1.7T parameters) and model performance (12.2% to 88.4% in terms of pass@1).
We further introduce each model as follows.
CodeGen-16B (Nijkamp et al., 2022) is an open-source LLM released by Salesforce. It employs a decoder-only architecture with rotary position embedding. The series of CodeGen models are trained on 217GB of Python code. We utilize the version with 16B parameters.
InCoder-1.3B (Fried et al., 2022) is another open-source LLM released by Meta AI. InCoder utilizes a new causal masking objective, which allows filling code blocks as well as standard left-to-right code generation. InCoder is trained on 159GB of open-source repositories with a permissive license from GitHub, GitLab, and StackOverflow. We adopt an InCoder variant with 1.3b parameters.
GPT-3.5 (cha, 2023) was released by OpenAI as the first model deployed for ChatGPT. It has been further refined for conversational tasks through the use of Reinforcement Learning with Human Feedback (RLHF). GPT-3.5 is trained with a massive number of crowd-sourced web texts up to September 2021.
GPT-4 (OpenAI, 2023) is a successor to GPT-3.5, developed and released by OpenAI, which is one of the most advanced language models available. Like GPT-3.5, GPT-4 has also undergone RLHF to enhance its conversational capabilities.
SantaCoder (Allal et al., 2023) is an open-sourced LLM released as a part of the BigCode project. It uses multi query attention and was trained with Python, Java, and Javascript subset of The Stack dataset with 1.1b parameters.
StarCoder (Li et al., 2023a) is another open-sourced LLM released as a part of the BigCode project. Compared with SantaCoder, it was trained on the complete The Stack dataset with 15.5b parameters.
| Model | Release | Size | Performance | ||
| Pass@1 | Pass@10 | Pass@100 | |||
| CodeGen-16B (Nijkamp et al., 2022) | Mar. 2022 | 16B | 32.9% | 56.0% | 81.5% |
| InCoder-1.3B (Fried et al., 2022) | Apr. 2022 | 1.3B | 12.2% | 15.9% | 25.2% |
| GPT-3.5 (cha, 2023) | Nov. 2022 | 175B | 73.2% | 88.6% | 94.0% |
| GPT-4 (OpenAI, 2023) | Mar. 2023 | 1.7T | 88.4% | — | — |
| SantaCoder (Allal et al., 2023) | Apr. 2023 | 1.1B | 14.6% | 29.3% | 45.9% |
| StarCoder (Li et al., 2023a) | May. 2023 | 15.5B | 34.1% | 56.7% | 84.2% |
2.3. Data Collection
In this study, we utilize the widely used HumanEval (Chen et al., 2021) to collect code generation errors made by LLMs. HumanEval includes 164 hand-written Python programming tasks, each accompanied by an average of 7.7 unit tests. These tasks involve language comprehension, reasoning, algorithms, and simple mathematics. For each task, We prompted each LLM with the original prompt from HumanEval and greedy decoding with the temperature set to 0. Then, we executed the test cases to identify incorrect solutions for the generated code solutions. We also performed a round of manual checks to find solutions that pass test cases but are not fully correct, since some tasks may not have sufficient test cases. In the end, we identified a total of 558 incorrect code solutions generated by the six models.
2.4. Manual Analysis Procedure
We followed the well-established open coding process (Strauss and Corbin, 2004; Antoine et al., 2021; Lazar et al., 2017; Braun and Clarke, 2006) to analyze the characteristics of the 558 incorrect code solutions and develop a taxonomy of code generation errors made by LLMs.
Open coding. From the 558 incorrect solutions, we first randomly sampled 160 of them as a starting point for analysis. The sample size is statistically significant, with a 95% confidence level and a 5% margin of error. Two authors independently labeled each incorrect solution and documented the errors in the code and their root causes. For some solutions, it is easy to identify the error and root cause by visually comparing the incorrect solution and the ground truth. Yet many solutions have subtle bugs or follow a different path to solve the task, e.g., using a lambda expression instead of a for loop to solve a problem. The authors had to copy the incorrect solution and their test cases to an IDE and performed step-by-step debugging to locate the errors.
The authors documented all error locations and root causes and discussed them with other authors after the initial coding. They refined some code labels in the meeting and came up with an initial codebook. At this stage, we found that code generation errors made by LLMs can be categorized along two dimensions based on their semantic and syntactic characteristics. Specifically, we identified seven common semantic characteristics and eight common syntactic characteristics of these errors.
Iterative refinement of the codebook. After obtaining the initial codebook, the two authors invited another two authors to iteratively improve the established codebook. The four authors were first assigned with a batch of 10 code generation errors. Subsequently, they independently labeled these code snippets from both semantic and syntactic perspectives. They were tasked with justifying their codes for every error and writing down the reasons. If a new error occurs that the codebook does not cover, an author needs to write a description of the error.
After the first round of labeling, we computed Fleiss’ Kappa (Fleiss, 1971) to determine the internal consistency among the authors. The initial scores were 0.37 and 0.32 for semantic characteristics and syntactic characteristics, respectively. To figure out where the disagreements were, the four authors met to discuss the differences between the justification of specific coding and exchange opinions about updating the codebook.
The four authors were then assigned a new batch of 10 errors and conducted another round of labeling following the aforementioned process. After the second round of labeling, the Fleiss’ Kappa scores were 0.68 and 0.69 for semantic characteristics and syntactic characteristics, respectively, indicating substantial agreement. They further conducted the third round of labeling with a new batch of 10 errors. After the third round of labeling, the Fleiss’ Kappa scores were 0.84 and 0.71 for semantic characteristics and syntactic characteristics, respectively, indicating substantial agreement. The final codebook was then established, which includes 13 common semantic characteristics and 14 common syntactic characteristics.
Annotating the remaining dataset. The second author further used the final codebook to label the remaining 390 incorrect code snippets. While labeling the remaining code snippets, the author found no new error characteristics. The final coding results were documented as a spreadsheet and provided in the supplementary material. The whole labeling process took about 268 person-hours.
3. Results
In the following subsections, we denote the 164 programming tasks in HumanEval (Chen et al., 2021) as Task 0-163. Due to the page limit, some of the code examples are omitted or simplified. We refer the readers to our website (https://llm-code-errors.cs.purdue.edu) for more details.
3.1. RQ1: What specific types of errors manifest in code generation across different LLMs?
Table 2 and Table 3.1.1 presents the finalized taxonomy of code generation errors made by LLMs. The taxonomy categorizes the incorrect code solutions based on their semantic and syntactic characteristics. The semantic characteristics highlight the high-level root causes of these code generation errors, while the syntactic characteristics indicate which parts of the code an error occurs in.
| Error Characteristics | Incorrect Example | Ground Truth | |
| Condition Error | Missing condition | ⬇ # [Task 151] CodeGen-16b def double_the_difference(lst): sum = 0 for i in lst: if i > 0 and i % 2 != 0: sum += i ** 2 return sum | ⬇ # [Task 151] Ground Truth def double_the_difference(lst): ans = 0 for num in lst: if num % 2 == 1 and num > 0 and "." not in str(num): ans += num ** 2 return ans |
| Incorrect condition | ⬇ # [Task 44] GPT-3.5 def change_base(x: int, base: int): result = "" while x > 0: remainder = x % base result = str(remainder) + result x = x // base return result | ⬇ # [Task 44] Ground Truth def change_base(x: int, base: int): if x == 0: return "0" ret = "" while x != 0: ret = str(x % base) + ret x //= base return ret | |
[…]
| Reference Error | Wrong method/variable | ⬇ # [Task 101] GPT-3.5 def words_string(s): words = s.split(", ") words = [word.split(" ") for word in words] return words | ⬇ # [Task 101] Ground Truth def words_string(s): words = (s.replace(",", " ")).split() return [word for word in words if word != ""] |
| Undefined name | ⬇ # [Task 83] CodeGen-16B def starts_one_ends(n): if n == 1: return 1 return starts_one_ends(n - 1) + ends_one_starts(n - 1) | ⬇ # [Task 83] Ground Truth def starts_one_ends(n): if n == 1: return 1 return 18 * 10 ** (n - 2) | |
| Operation/ Calculation Error | Incorrect arithmetic operation | ⬇ # [Task 44] InCoder-1B def change_base(x: int, base: int): digits = [] while x: digits.append(str(x % base)) x /= base return ’’.join(reversed(digits)) | ⬇ # [Task 44] Ground Truth def change_base(x: int, base: int): if x == 0: return "0" ret = "" while x != 0: ret = str(x % base) + ret x //= base return ret |
| Incorrect comparison operation | |||
| Garbage Code | Only comments | ⬇ # [Task 152] InCoder-1B def compare(game,guess): # Regenerate the task description. | ⬇ # [Task 152] Ground Truth def compare(game,guess): return [abs(game[i] - guess[i]) for i in range(len(game))] |
| Meaningless code snippet | ⬇ # [Task 159] CodeGen-16B def eat(number, need, remaining): pass | ⬇ # [Task 159] Ground Truth def eat(number, need, remaining): if need <= remaining: return [number + need, remaining - need] else: return [number + remaining, 0] | |
| Wrong (logical) direction | ⬇ # [Task 20] InCoder-1.3B def find_closest_elements(numbers): closest_to_one, closest_to_two = numbers[0], numbers[1] for number in numbers: if number < closest_to_one: closest_to_one = number if number > closest_to_two: cloeset_to_two = number return closest_to_one, closest_to_two | ⬇ # [Task 20] InCoder-1.3B def find_closest_elements(numbers): min_diff, min_pair = float("inf"), None for l, r in zip(numbers[:-1], numbers[1:]): diff = r - l if diff < min_diff: min_diff = diff min_pair = (l, r) return min_pair | |
| Incomplete Code/ Missing Steps | Missing one step | ⬇ # [Task 16] InCoder-1B def count_distinct_characters(string: str) -> int: return len(set(string)) | ⬇ # [Task 16] Ground Truth def count_distinct_characters(string: str) -> int: return len(set(string.lower())) |
| Missing multiple steps | |||
| Memory Error | Infinite loop | ⬇ # [Task 100] CodeGen-16b def make_a_pile(n: int): if n % 2 == 0: return [n] + make_a_pile(n + 2) else: return [n] + make_a_pile(n + 1) | ⬇ # [Task 100] Ground Truth def make_a_pile(n: int): ans, num = [], n for _ in range(n): ans.append(num) num += 2 return ans |
3.1.1. Semantic Characteristics.
We identified the following 7 types and 13 sub-types of common semantic characteristics in the code generation errors.
Condition Error includes missing condition and incorrect condition. Missing condition is when necessary conditions are omitted, while incorrect condition is when conditions are incorrectly formulated, leading to errors.
Constant Value Error is an error that occurs when an incorrect constant value is set. An incorrect constant value can occur in function arguments, assignments, or other parts of the code.
Reference Error involves incorrect references to variables or functions, which includes the usage of an incorrect function or variable that does not match the requirement (wrong method/variable) and reference to a variable or method name that has not been defined (undefined name).
Operation/Calculation Error indicates the mistakes in mathematical or logical operations.
Garbage Code is defined as unnecessary parts of the code that do not contribute to the intended functionality. It can occur in several forms: a meaningless code snippet, where the code, though syntactically correct, is irrelevant to the assigned task; only comments, where the code consists exclusively of comments without any executable statements to address the task; or wrong (logical) direction, where the code significantly deviates from the intended task logic and expected outcomes.
Incomplete Code/Missing Steps indicates the absence of crucial steps to achieve the task.
Memory Error includes infinite loop, which is a loop or recursion that never terminates.
Comparison between LLMs. Fig. 1 shows the distribution of the 13 semantic characteristics for each LLM. On the one hand, we find that several characteristics are frequently shared among all LLMs, such as incorrect condition and wrong (logical) direction. This implies that all LLMs struggle with certain kinds of task requirements such as handling complex logic conditions, regardless of their model size and capability. For instance, CodeGen-16B generated an incorrect condition in Example 1 (Task 146). In this example, the task requires the program to check if the first and last digits of a number are odd. However, CodeGen-16B generated an incorrect condition to check if the number itself is odd.
On the other hand, small models such as InCoder and CodeGen are more likely to generate meaningless code and code that miss multiple steps, while larger models such as GPT-3.5 and GPT-4 tend to make more constant value errors and arithmetic operation errors. Notably, incorrect code generated by GPT-4 only exhibited 7 of the 13 semantic characteristics, while incorrect code generated by smaller models exhibited all sorts of errors. One plausible explanation is that GPT-3.5 and GPT-4 are much larger and thus are better at interpreting task descriptions. For instance, both CodeGen-16B and InCoder-1.3B missed the subtle requirement in Example LABEL:exp3 (Task 85) that “the elements must be even numbers”, resulting in missing steps within the generated code. However, GPT-3.5 and GPT-4 were able to generate the correct solution successfully.
CodeGen-16B
InCoder-1.3B
GPT-3.5
GPT-4
Loop ErrorFor error
While error
Return Error Incorrect return value
Method Call Error Incorrect function name
Incorrect function arguments
Incorrect method call target
Assignment Error Incorrect constant
Incorrect arithmetic
Incorrect variable name
Incorrect comparison
Import Error Import error
Code Block Error Incorrect code block
Missing code block
Finding 1: The most common semantic characteristics among six LLMs are wrong (logical) direction and incorrect condition. Compared with the other four LLMs, GPT-3.5 and GPT-4 do not often generate code with missing steps.
Furthermore, we found that even for the same failure task, failed code generated by different LLMs could have different semantic characteristics. Example LABEL:exp-failure-case shows an example of three LLMs that made different semantic characteristics. In this example, CodeGen-16B misses a few steps of handling the “unit digits” (missing multiple steps), while InCoder-1.3B completely ignored the task (wrong (logical) direction). GPT-3.5 only made a small mistake with parentheses (incorrect arithmetic operation). One takeaway from such observation is that ensemble of different LLMs might improve the code generation success rate.
CodeGen-16B
GPT-3.5
InCoder-1.3B
Also, 37.5% are only comments, defined as the generated code consists solely of comments and lacks executable statements to solve the given task. Besides, 25.0% of them fall into wrong (logical) direction, indicating a significant misinterpretation of the task requirements. In these cases, the LLM generates code that not only fails to solve the problem but also produces solutions that contradict the intended outcome. For instance, in Task 115, GPT-3.5 wrongly assumes that water from different wells can be combined into a single bucket operation, whereas each well should be considered separately according to its capacity constraints.
3.2. RQ4: How does code length relate to the types of LLMs’ code generation errors?
We did not find any statistical evidence that the generated code length is related to a specific task’s success. As shown in Fig. 2, in most cases, the passed tasks’ code length is similar to the failed tasks’ code length on all six LLMs. Mann-Whitney -test (Mann and Whitney, 1947) results further suggest that the mean differences are not statistically significant in all cases (). Finding 3: There is no statistical evidence indicating that there is a difference between the passed tasks’ code length and the failed tasks’ code length.
Though the average lengths of the code generated by CodeGen-16B, InCoder-1.3B, GPT-3.5, GPT-4, SantaCoder, and StarCoder (29.3, 32.3, 25.2, 29.6, 26.5, and 15.4 words, respectively) are short, we noticed a small portion of the generated code with a large length (i.e., longer than 100 words). Therefore, we further investigated these long code snippets generated by LLMs. In particular, we found that out of 36 code snippets with a length longer than 100 words, only 4 of them were correct. These results indicate that when the generated code is extremely long, there is a small chance that the code is correct. We further dug into those failure cases. Specifically, we found that 28 out of the 32 failed code snippets (87.5%) are Garbage Code, where 22 are Meaningless code snippet, 4 are Only comments, and only 2 are Wrong (logical) direction. Regarding Meaningless code snippet, we manually checked the 22 code snippets and found that all of them encountered the same problem: repeatedly generating similar lines/blocks of code. For instance, Task 6 requires the program to output the deepest level of nested parentheses. However, SantaCoder repeatedly generated the code to replace the parentheses with empty strings.
3.3. RQ5: Does partially failed code exhibit different characteristics compared with fully failed code?
Since HumanEval (Chen et al., 2021) provides test cases for each task, we are interested in whether a completely failed task (all test cases failed) would yield different error characteristic patterns compared with a partially failed task (a subset of test cases failed). Overall, we find that there are 244 completely failed tasks and 314 partially failed tasks among 558 incorrect code snippets. Fig. 3 shows the distribution of task failure rates for incorrect solutions with different semantic characteristics. Only comments and meaningless code are the top two characteristics that lead to complete failures. Despite a small mistake, the majority of code with undefined names also leads to complete test failures due to compile-time or runtime crashes. Surprisingly, we found that while some error characteristics, such as missing multiple steps, sound severe by definition, they do not often lead to complete failures. After digging into the concrete instances, we noticed that this is because LLMs do not completely misunderstand the task description and the generated code can still pass some weak test cases. For instance, Task 125 requires the program to split on commas if there is no space in the string. If there is also no comma in the string, the program should perform other operations. However, the code generated by CodeGen-16B only split the input string into spaces while missing the remaining steps. As a result, it can only pass the test cases that include spaces, leading to a partially failed task.
Another surprising finding is that code with the wrong direction can pass some test cases accidentally. For example, in Task 75, the code generated by InCoder-1.3B does not follow the task instructions. However, it can pass a few test cases, such as 5, 10, and 30. One possible reason is that since the prompt includes test cases, LLMs may have memorized a superficial correlation between test cases and some other irrelevant solutions that pass those test cases.
In terms of syntactic characteristics (Fig. 4), we observed an obvious difference on if error, where the majority of code with this error led to partial failure (84% v.s. 16%). This is because when encountering if error, LLMs were more likely to only misinterpret a specific condition requirement while generating correct code for the other parts. For instance, in Task 0, SantaCoder only considered the adjunct elements in the given list numbers. As a result, it failed to pass the test case where the two closet elements were not adjunct.
4. Implications and Future Work
In this section, we discuss several significant implications of our study for code generation by LLMs and research opportunities to improve the quality and reliability of code LLMs.
Implication 1: Prompt engineering for better code generation. Our Finding 3.1.1 reveals that when the input prompt is too long, the LLMs may fail to understand the given complex task requirements, leading to Garbage Code generation. By contrast, a short and concise task description usually leads to better code generation. Therefore, optimizing long and complex task requirements into short and concise prompts could be another research opportunity. This could be highly related to the field of prompt engineering. Recent studies from other domains (e.g., computer vision and natural language processing) have shown the promise of prompt engineering techniques such as chain-of-thought prompting (Zhao et al., 2021; Wei et al., 2022), automated prompt refinement (Pavlichenko and Ustalov, 2023), and interactive prompt engineering tools (Strobelt et al., 2022; Feng et al., 2023). Simultaneously, software engineering researchers have also been working on prompt tuning and prompt design for code generation (Wang et al., 2022b). Our work further offers practical insights into how the length of the prompt could affect the success of LLMs’ code generation. Therefore, in future work, researchers may work on investigating to what extent the existing prompt engineering techniques can facilitate code generation tasks or propose new techniques to optimize prompts.
Implication 2: Developing fault localization techniques for LLMs’ code generation. Our Finding 3.1.1 and Finding LABEL:finding:2 reveal that non-trivial efforts may be required to repair incorrect code snippets generated by LLMs. This could be largely attributed to the low similarity between the incorrect code snippets and the correct ground truth code, as well as the diversity of error characteristics that LLMs may produce. As a result, identifying the root causes of these incorrect code snippets becomes challenging. Without precise fault localization, the effectiveness of automated program repair techniques could also be significantly affected (Wen et al., 2019). Although there has been a large body of work about fault localization approaches (especially spectrum-based ones) in SE (Wen et al., 2019; Abreu et al., 2009), we argue that two obvious challenges still exist for fault localization in LLMs’ generated code. First, spectrum-based fault localization usually targets generating a fault space at statement granularity (Wen et al., 2019). However, our study results reveal that LLMs tend to produce a significant number of errors at the block level (e.g., incorrect code block and missing code block). In these cases, statement-level fault localization may be ineffective. Second, existing fault localization methods may not distinguish different types of error characteristics in an incorrect code snippet. Our study results show that the incorrect code snippets generated by LLMs could include not only simple characteristics (e.g., incorrect condition, which may be addressed by existing APR techniques) but also complex ones such as wrong (logical) direction. These issues are usually related to the LLMs’ misinterpretation of task requirements. Without accurately identifying these types of errors, fixing them could be even more challenging. Therefore, in future work, researchers may consider developing techniques to automatically identify the types of semantic and syntactic error characteristics in LLM-generated code, considering the unique patterns and complexities introduced by LLMs.
Implication 3: Developing error repair techniques for LLMs’ code generation. Our study discovers that some of the syntactic characteristics may be easy to fix, such as if error and incorrect function arguments. To repair these one-line errors, practitioners could leverage existing techniques for automated program repair (APR) (Tan and Roychoudhury, 2015; Long and Rinard, 2016; Mechtaev et al., 2016; Saha et al., 2017; Xiong et al., 2017; Ghanbari et al., 2019; Lutellier et al., 2020; Jiang et al., 2018; Wen et al., 2018; Liu et al., 2019; Li et al., 2022b; Ye et al., 2022; Fan et al., 2023; Xia and Zhang, 2022; Jiang et al., 2023a, b). However, applying APR techniques could still be challenging in some cases. For instance, if error could originate from the miss of specific task requirements by LLMs (Task 0 in Sec. 3.3), where APR techniques may not work. Moreover, our findings reveal that a larger number of syntactic characteristics made by code LLMs could cover multiple lines and statements (e.g., incorrect code block, missing code block), which also result in the low similarity between the incorrect code snippets and the correct ground truth code. Such errors usually go beyond simple fixes and require substantial restructuring and logical corrections. It would be worth investigating to what extent the novel LLM-based APR methods could resolve these issues (Fan et al., 2023), as well as developing new techniques to repair block-level code errors.
5. Related Work
5.1. LLM-based Code Generation
In recent years, Large Language Models (LLMs) have not only demonstrated great performance in various natural language processing tasks (Dong et al., 2019; cha, 2023), but also shown great potential in code-related tasks, such as generation (Chen et al., 2021; Fried et al., 2022; Nijkamp et al., 2022; Wang et al., 2023; Zhu et al., 2024), summarization (Ahmad et al., 2020; Niu et al., 2022; Cai et al., 2024), understanding (Wang et al., 2022a; Ding et al., 2024b), search (Feng et al., 2020; Zhang et al., 2022b; Jain et al., 2023) and translation (Roziere et al., 2020). As such, several LLMs specifically trained or fine-tuned on code data have been introduced, e.g., AlphaCode (Li et al., 2022a), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023a), CodeGen (Nijkamp et al., 2022) and Qwen (Bai et al., 2023). A notable example is CodeLlama (Roziere et al., 2023), an adaptation of Llama-2 (Touvron et al., 2023). To enhance its code understanding and generation, CodeLlama introduced three significant modifications: (1) the collection of higher-quality code data for training, (2) the implementation of Fill-in-the-Middle (FIM) task (Fried et al., 2022) and (3) the adoption of instruction fine-tuning to better respond to code-related queries. Additionally, CodeLlama employed the self-instruct technique, where models generate programming questions and answers that are subsequently used for fine-tuning. Follow-up work can then be roughly categorized into three groups: (1) those that focus more on collecting more diverse, high-quality code-related training data, (2) those that seek better instruction fine-tuning techniques, and (3) those that utilize existing LLMs and devise better prompting strategies (Zhang et al., 2023; Bairi et al., 2024; Gao et al., 2023; Chen et al., 2023; Olausson et al., 2023; Li et al., 2023b; Ding et al., 2024a). For the first line of work, StarCoder 2 (Lozhkov et al., 2024) built the Stack v2 dataset, which is 4x larger than its predecessor, with up to 4.3 trillion tokens from more diverse sources such as GitHub pull requests, Kaggle notebooks, and code documentation. Additionally, DeepSeek-coder (Guo et al., 2024) introduced an approach for analyzing dependencies between code files using topological sorting, aggregating linked code lines into a single training sample, enabling LLMs to understand repository-level code bases better. For the second line of work, a representative example is WizardCoder (Luo et al., 2024), which implemented an “instruction evolution” strategy, where LLMs are iteratively guided to generate increasingly complex instructions covering a broader range of topics and skills. Later, MagicCoder (Wei et al., 2023) introduced OSS-Instruct, a dataset where GPT-3.5 Turbo is prompted to create coding problems and their solutions based on collected code snippets. OSS-Instruct consists of 75k fine-tuning instances, and the resulting MagicCoder, based on Code Llama, has demonstrated superior performance over GPT-3.5 on the HumanEval benchmark. The third line of the work focuses on designing more complex prompting strategies based on existing pret-trained LLMs. For example, Self-Debugging (Chen et al., 2024) enables LLMs to debug their own code by prompting with execution results and instructions for code explanations, demonstrating better accuracy than GPT-3.5 and GPT-4 on the MBPP dataset (Austin et al., 2021). Self-Refine (Madaan et al., 2024) iteratively enhances code generation by instructing the LLM to provide feedback on its outputs and refine them, achieving up to 13% improvement on CodeX (Chen et al., 2021). Cycle (Ding et al., 2024a) also leverages a similar idea of self-refinement based on the execution results of test suites. Different from Self-Debugging and Self-Refine, the Cycle includes a fine-tuning stage to enhance LLMs’ abilities to refine faulty programs. As a result, Cycle outperforms models with 3 times more parameters in self-refinement. CodeT adopts a test-driven prompting method (Chen et al., 2023). Rather than relying on existing test suites, CodeT instructs LLMs to simultaneously generate multiple code snippets and their corresponding tests, selecting the final output based on the highest consensus. SKCoder (Li et al., 2023b) utilizes a retriever-based pipeline to extract code sketches from existing databases. Distinct from these efforts that aim to train better LLMs, our research focuses on analyzing the errors LLMs produce. We aim to develop a taxonomy of these errors, with the hope that our findings will assist the development of new methods in the direction and contribute to better LLM-enabled intelligent software engineering.
5.2. Quality of LLM-generated Code
Evaluating the quality of LLM-generated code can reveal the current approach’s shortcomings and guide future improvement. While prior research has delved into various facets such as robustness (Arakelyan et al., 2023; Liu et al., 2024b), security (Pearce et al., 2022), and usability (Barke et al., 2023; Al Madi, 2022), studies specifically examining the correctness of LLM-generated code (Chen et al., 2021; Nguyen and Nadi, 2022; Jesse et al., 2023; Liu et al., 2023, 2024a) are of particular relevance to our work. Liu et al. (Liu et al., 2023) evaluated the quality of code generated by ChatGPT, addressing various factors, including compilation and runtime errors, output correctness, coding style, maintainability, and performance. Their work, however, focused more on concrete error types. By contrast, our work focuses on code generation. We investigate a broader range of LLMs and examine the error characteristics’ correlations with prompt complexity, code similarity, etc. Pan et al. (Pan et al., 2024) introduced a taxonomy centered on code translation bugs. These bug types include code translate-specific problems (e.g., semantics differences between different programming languages) and logic and data-related bugs (e.g., the inclusion of logic not in source code and mismatched output data structure), which share some similarities with us. However, there is a key distinction in the origin of errors where Pan et al.’s work focuses on code-to-code translation, while our focus is on the issues in LLMs’ code generation from natural language specifications. Finally, Liu et al. (Liu et al., 2024a) conducted a study of the quality of code generated by ChatGPT, assessing their correctness, understandability, and security. In their examination of code correctness, they primarily focused on compile errors and runtime crashes. Our research differs from theirs in two key respects: (1) Our study subjects are more diverse with both open-source and closed-source models; (2) our taxonomy also considers behavior deviations informed by test failures in addition to compiling errors and crashes. Based on the findings, we further provide actionable suggestions and implications for future enhancements in LLM-enabled code generation.
5.3. Taxonomy on Software Defects
In software engineering, collecting, classifying, and analyzing software defects has been recognized as an important research topic for improving software quality. Building a systematic taxonomy of software bugs can help stakeholders understand the pitfalls of target systems and guide developers for better development practices. One of the early attempts along the direction was the orthogonal defect classification (ODC) proposed by IBM Research (Chillarege et al., 1991, 1992). Since then, numerous endeavors have been made to construct defect taxonomies from various angles. These include investigations targeting different programming languages, such as C (Le Goues et al., 2015; Tan et al., 2017), Javascript (Hanam et al., 2016; Gyimesi et al., 2019), Java (Pan et al., 2009), Python (Zhang et al., 2022a), Julia (Rahman et al., 2023), and multi-language programs (Lin et al., 2017; Kochhar et al., 2016). Others delve into different applications, such as Infrastructure as Code (IaC) (Rahman et al., 2020) and deep learning stacks (Humbatova et al., 2020). Different from these attempts, our work mainly focuses on building a systematic taxonomy for LLM-generated code.
6. Threats to Validity
Threats to Internal Validity. One potential threat comes from our manual analysis process, where labelers may have different opinions and sometimes may make mistakes. To mitigate this, four of the authors first performed open-coding and iteratively refined our codebook until a substantial agreement was achieved before one of the authors labeled the complete dataset. Our final Fleiss’ Kappa regarding the semantic and syntactic characteristics are 0.84 and 0.71, respectively, indicating substantial agreement.
Threats to External Validity. One potential threat lies in the choice of dataset. Considering the extensive labeling effort (e.g., running the code in an IDE and performing step-by-step debugging to locate the root cause), we have only labeled one dataset. In future work, one may consider labeling more datasets to confirm the generalizability of our findings. Nevertheless, given the size of our labeled dataset (558 code snippets), we believe our findings can be generalized to other similar datasets, such as MBPP and APPS. Besides, our study only covers Python programs, which might not generalize to programming languages that are very different from Python (e.g., PHP and Rust). We chose Python because it is one of the most popular object-oriented programming languages. In future work, we plan to expand our study with programming tasks from other languages.
Threats to Construct Validity. In RQ3, we use the prompt length to estimate the prompt complexity. Although prompt length may not be the best metric for such estimation, we believe this is a pragmatic choice since there is no commonly used metric for prompt complexity. In future work, researchers may consider designing new metrics to estimate the prompt complexity and investigate its correlation with LLM’s code generation capabilities.
7. Conclusion
In this paper, we present an empirical study on code generation errors made by large language models. We first derived a taxonomy of LLMs’ code generation errors based on six popular LLMs’ failure cases on the HumanEval dataset (Chen et al., 2021) through open coding and thematic analysis. We labeled a total of 558 errors committed by these code LLMs according to the established taxonomy. We found that these LLMs exhibited different distributions of semantic and syntactic error characteristics. We further analyzed these error characteristics from various perspectives, such as prompt complexity and test-pass rate. At the end of the paper, we discuss the implications of our study and propose future research opportunities for improving the quality and reliability of code LLMs.