### v2

# Where Do Large Language Models Fail When Generating Code?

Abstract.
Large Language Models (LLMs) have shown great potential in code generation. However, current LLMs still cannot reliably generate correct code. Moreover, it is unclear what kinds of code generation errors LLMs can make. To address this, we conducted an empirical study to analyze incorrect code snippets generated by six popular LLMs on the HumanEval dataset. We analyzed these errors alongside two dimensions of error characteristics—semantic characteristics and syntactic characteristics—to derive a comprehensive code generation error taxonomy for LLMs through open coding and thematic analysis. We then labeled all 557 incorrect code snippets based on this taxonomy. Our results showed that the six LLMs exhibited similar distributions of syntactic characteristics while different distributions of semantic characteristics. Furthermore, we analyzed the correlation between different error characteristics and factors such as task complexity, code length, and test-pass rate. Finally, we highlight the challenges that LLMs may encounter when generating code and propose implications for future research on reliable code generation with LLMs.
1. Introduction
Automatically generating code from natural language has been a long-term pursuit across multiple research communities. Recent advances in Large Language Models (LLMs) have led to rapid, unprecedented improvements on this task (Nijkamp et al., 2022; Fried et al., 2022; Li et al., 2023a; OpenAI, 2023; cha, 2023). Despite this great progress, LLMs still cannot reliably generate correct code for many tasks. Currently, there is a lack of deep understanding of the cases where LLMs fail. Specifically, it remains unclear what types of code generation errors an LLM typically produces and whether different LLMs make similar errors. Answering these questions would help researchers gain insights into the limitations of existing models and identify opportunities for model improvement.
To bridge this knowledge gap, we conducted an in-depth analysis of code generation errors made by LLMs. We focused on six popular LLMs: CodeGen-16B (Nijkamp et al., 2022), InCoder-1.3B (Fried et al., 2022), GPT-3.5 (cha, 2023), GPT-4 (OpenAI, 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023a). These models produced 557 incorrect code solutions on the 164 tasks from the HumanEval dataset (Chen et al., 2021). Four of the authors worked together to locate the erroneous parts of these incorrect solutions and manually fix them. Specifically, for some tasks, LLMs may propose an alternative solution that differs from the ground truth solution in HumanEval, e.g., using a lambda expression to process a sequence of data instead of using a loop. To avoid overfitting the ground truth, the authors manually locate and fix errors following the problem-solving direction of the LLM instead of simply comparing the LLM-generated code with the ground truth.
We performed multiple rounds of open coding and iterative refinement to analyze the characteristics of the located errors. Specifically, we analyzed these errors alongside two dimensions—the semantic characteristics and syntactic characteristics of these errors:
Semantic characteristics can help identify the high-level root causes of these code generation errors. Representative semantic characteristics include missing condition, wrong (logical) direction, incorrect condition, etc. Analyzing these semantic characteristics across various LLMs can help understand the limitations of current LLMs in interpreting task requirements and generating semantically correct programs.
Syntactic characteristics can help localize where the error occurs in an incorrect code solution. Representative syntactic characteristics include incorrect code blocks, incorrect function arguments, etc. Understanding these characteristics allows for a better assessment of current LLMs’ abilities to generate different kinds of code constructs. It can also help inform the design of new techniques for localizing and repairing code generation errors made by LLMs.
Our analysis shows that the majority of code generation errors involve multiple lines of code, rather than simple errors. These errors often require substantial code restructuring and repair rather than simple fixes. Furthermore, while the overall distribution of the syntactic characteristics of these errors (i.e., error locations) is similar across different LLMs, the semantic characteristics of the errors (i.e., root causes) vary significantly for different LLMs even for the same task. Interestingly, most of the incorrect code solutions are compilable and runnable without any compilation errors. Thus, we cannot easily capture these errors via compiler check. Careful code review and high-quality test cases are necessary to capture these errors. This also implies that modern LLMs have adequately learned the syntax rules of programming languages, but struggle with understanding intricacies in natural language task descriptions and generating delicate code with sophisticated logic.
In summary, this paper makes the following contributions:
We established a taxonomy of both syntactic and semantic characteristics of code generation errors through open coding and thematic analysis.
We analyzed the similarities and differences in the errors made by different LLMs, as well as the bug-fixing effort, the impact of task complexity and code length, and the correlation between test pass rates and different kinds of errors.
We discussed the implications and future opportunities for improving LLMs for code generation.
We developed an interactive data analysis website to help researchers and developers examine and explore code generation errors in different categories. The website is available at https://llm-code-errors.cs.purdue.edu.
2. Methodology
In this section, we first introduce the research questions of this study, followed by the descriptions of the dataset and LLMs used in our study and the procedure of our manual analysis.
2.1. Research Questions
This study investigates the following research questions.
RQ1: What kinds of code generation errors do different LLMs make? This question aims to uncover the common characteristics and distinctions of code generation errors made by different LLMs. This can help us understand whether it is feasible to develop generic methods to improve LLMs or whether these models need specialized treatment.
RQ2: How much effort is needed to fix code generation errors? In practice, it is unrealistic to expect LLMs to generate fully correct code for every possible scenario. Existing studies show that some incorrect code can still serve as a useful starting point for developers (Vaithilingam et al., 2022; Barke et al., 2023). Thus, it is important to understand what efforts are needed to fix the incorrect solutions and whether it is possible to automate the repair. This question aims to fill this knowledge gap.
RQ3: How does the task complexity affect an LLM’s code generation? Intuitively, complex tasks are more challenging to solve than simple tasks. Yet it is unclear whether different LLMs exhibit different code generation capabilities when solving tasks of different complexity levels. Specifically, it would be useful to find out whether there is an upper bound on the complexity of tasks that LLMs caneloquently solve, which can then be used to guide or estimate the effort required for code review, testing, and repair.
RQ4: How does code length relate to the types of LLMs’ code generation errors? Different from RQ3, this question aims to investigate if the generated code length affects the correctness of a code snippet generated by LLMs. We are particularly interested in those lengthy code solutions and their error characteristics.
RQ5: Does partially failed code exhibit different characteristics compared with fully failed code? This question explores the distinctions between code that fails a subset of test cases and code that fails all test cases. It can offer insights into the specific challenges faced in achieving full correctness.
2.2. Code Generation LLMs
In this study, we focus on six representative code generation LLMs: CodeGen-16B (Nijkamp et al., 2022), InCoder-1.3B (Fried et al., 2022), GPT-3.5 (cha, 2023), GPT-4 (OpenAI, 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023a). As shown in Table 1, these models cover a wide range of model sizes (1B to 1.7T parameters) and model performance (12.2% to 88.4% Pass@1). As GPT-3.5 and GPT-4 are constantly evolving, we used GPT-3.5-Turbo-0301 and GPT-4-0314, which were the two most recent model checkpoints at the time of our analysis.
We further introduce each model as follows.
CodeGen-16B (Nijkamp et al., 2022) is an open-source LLM released by Salesforce. It employs a decoder-only architecture with rotary position embedding. The series of CodeGen models are trained on 217GB of Python code. We utilize the version with 16B parameters.
InCoder-1.3B (Fried et al., 2022) is another open-source LLM released by Meta AI. InCoder utilizes a new causal masking objective, which allows filling code blocks as well as standard left-to-right code generation. InCoder is trained on 159GB of open-source repositories with a permissive license from GitHub, GitLab, and StackOverflow. We adopt an InCoder variant with 1.3b parameters.
GPT-3.5 (cha, 2023) was released by OpenAI as the first model deployed for ChatGPT. It has been further refined for conversational tasks through the use of Reinforcement Learning with Human Feedback (RLHF). GPT-3.5 is trained with a massive number of crowd-sourced web texts up to September 2021.
GPT-4 (OpenAI, 2023) is a successor to GPT-3.5, developed and released by OpenAI, which is one of the most advanced language models available. Like GPT-3.5, GPT-4 has also undergone RLHF to enhance its conversational capabilities.
SantaCoder (Allal et al., 2023) is an open-sourced LLM released as a part of the BigCode project. It uses multi query attention and was trained with Python, Java, and Javascript subset of The Stack dataset with 1.1b parameters.
StarCoder (Li et al., 2023a) is another open-sourced LLM released as a part of the BigCode project. Compared with SantaCoder, it was trained on the complete The Stack dataset with 15.5b parameters.
| Model | Release | Size | Performance | |
| Pass@1 | Incorrect Solutions | |||
| CodeGen-16B (Nijkamp et al., 2022) | Mar. 2022 | 16B | 32.9% | 110 |
| InCoder-1.3B (Fried et al., 2022) | Apr. 2022 | 1.3B | 12.2% | 144 |
| GPT-3.5 (cha, 2023) | Nov. 2022 | 175B | 73.2% | 42 |
| GPT-4 (OpenAI, 2023) | Mar. 2023 | 1.7T | 89.0% | 18 |
| SantaCoder (Allal et al., 2023) | Apr. 2023 | 1.1B | 14.6% | 139 |
| StarCoder (Li et al., 2023a) | May. 2023 | 15.5B | 34.1% | 104 |
2.3. Collection of Incorrect Code Solutions
In this study, we utilize the widely used HumanEval benchmark (Chen et al., 2021) to collect code generation errors made by LLMs. HumanEval includes 164 hand-written Python programming tasks, each accompanied by an average of 7.7 unit tests. These tasks involve language comprehension, reasoning, algorithms, and simple mathematics. For each task, we prompted each LLM with the original prompt from HumanEval and greedy decoding with the temperature set to 0. Then, we executed the test cases to identify incorrect solutions. We also performed a round of manual checks to find solutions that pass test cases but are not fully correct since some tasks may not have sufficient test cases. In the end, we identified a total of 557 incorrect code solutions generated by the six models. Table 1 shows the distribution of these incorrect solutions.
2.4. Manual Analysis of Incorrect Code Solutions
We followed the well-established open coding process (Strauss and Corbin, 2004; Antoine et al., 2021; Lazar et al., 2017; Braun and Clarke, 2006) to analyze the characteristics of the 557 incorrect code solutions and develop a taxonomy of code generation errors made by LLMs.
Open coding. From the 557 incorrect solutions, we first randomly sampled 160 of them as a starting point for analysis. The sample size is statistically significant, with a 90% confidence level and a 5% margin of error. Two authors independently identified the erroneous parts of each incorrect solution and documented the root causes of the errors. Since LLMs may generate alternative solutions compared with the ground-truth solution from HumanEval, we chose to manually debug the incorrect solution rather than simply comparing it with ground truth. Specifically, the two authors executed the failed test cases and performed step-by-step debugging to locate the errors and identify their root causes.
The authors documented all error locations and root causes and discussed them with other authors after the initial coding. They refined code labels in the meeting and came up with an initial codebook. At this stage, we found that code generation errors made by LLMs can be categorized along two dimensions based on their semantic and syntactic characteristics. Specifically, the initial codebook includes seven semantic characteristics and eight syntactic characteristics of code generation errors.
Iterative refinement of the codebook. After obtaining the initial codebook, the two authors invited another two authors to iteratively improve the established codebook. The four authors first independently analyzed 10 incorrect code snippets following the same procedure described above and labeled the error characteristics based on the initial codebook. If a new characteristic was identified, an author created a new label to describe the characteristic.
After the first round of labeling, we computed Fleiss’ Kappa (Fleiss, 1971) to measure the inter-rater agreement. We used Fleiss’ Kappa instead of Cohen’s Kappa, since we had more than two labelers and more than two labels. The initial scores were 0.37 and 0.32 for semantic characteristics and syntactic characteristics, respectively, indicating fair agreement (Landis and Koch, 1977). To figure out where the disagreements were, the four authors met to discuss the differences between the justification of specific coding and exchange opinions about updating the codebook. They found that the low agreement was due to missing error characteristics in the initial codebook.
The four authors then refined the codebook with newly identified error characteristics and labeled another batch of 10 incorrect solutions. The Fleiss’ Kappa scores of this round of labeling were 0.68 and 0.69 for semantic characteristics and syntactic characteristics, respectively, indicating substantial agreement (Landis and Koch, 1977). The authors further discussed the disagreements and refined the codebook. Then, they conducted the third round of labeling with a new batch of 10 errors. The authors did not find any new error characteristics at this round of labeling, and the Fleiss’ Kappa scores also increased to 0.84 and 0.71. At this point, the authors believed that the codebook was comprehensive enough. The final codebook includes 13 semantic characteristics and 14 syntactic characteristics.
Analyzing the remaining dataset. The two authors used the final codebook to label the remaining incorrect code snippets. The final Fleiss’ Kappa scores were 0.91 and 0.91 for semantic and syntactic characteristics, respectively, indicating perfect agreement (Landis and Koch, 1977). They also found no new error characteristics. The final coding results were documented in a spreadsheet and provided on our website at https://llm-code-errors.cs.purdue.edu. The whole labeling process took about 328 person-hours.
2.5. Analysis of Repair Effort
To investigate the repair effort (RQ2), we employ two different metrics to measure the similarity between incorrect model-generated code and the corresponding correct solution. First, we used Jaccard similarity (Jaccard, 1901), which treats code as a set of tokens and measures code similarity by computing the token overlap between two code snippets. Second, we used Levenshtein distance (Levenshtein et al., 1966) to compute the minimum number of edits (i.e., insertions, deletions, or substitutions) required to change an incorrect solution to the correct solution. Note that for some tasks, an LLM may propose an alternative solution with errors compared with the ground-truth solution in HumanEval. In the previous labeling process, we identified 17 incorrect solutions where the LLM proposed an alternative way to solve the task but did not correctly solve it. In such cases, it is unfair to directly compare the incorrect code with the ground truth. To address this issue, one author manually solved the task following the LLM’s solution and computed the metrics by comparing the incorrect solution with the alternative, correct solution.
3. Results
In the following subsections, we denote the 164 programming tasks in HumanEval (Chen et al., 2021) as Task 0-163. Due to the page limit, some code examples are simplified. We refer the readers to our website (https://llm-code-errors.cs.purdue.edu) for more details.
3.1. RQ1: What kinds of code generation errors do different LLMs make?
Table 2 and Table 3.1.1 presents the finalized taxonomy of code generation errors made by LLMs. The taxonomy categorizes code generation errors based on their semantic characteristics (i.e., root causes) and syntactic characteristics (i.e., error locations). In total, there are 13 semantic characteristics in 7 categories and 14 syntactic characteristics in 7 categories. We elaborate on each of them below.
| Error Characteristics | Incorrect Example | Correct Solution | |
| Condition Error | Missing condition | ⬇ # [Task 151] CodeGen-16b def double_the_difference(lst): sum = 0 for i in lst: if i > 0 and i % 2 != 0: sum += i ** 2 return sum | ⬇ # [Task 151] Ground Truth def double_the_difference(lst): ans = 0 for num in lst: if num % 2 == 1 and num > 0 and "." not in str(num): ans += num ** 2 return ans |
| Incorrect condition | ⬇ # [Task 44] GPT-3.5 def change_base(x, base): result = "" while x > 0: remainder = x % base result = str(remainder)+result x = x // base return result | ⬇ # [Task 44] Ground Truth def change_base(x, base): if x == 0: return "0" ret = "" while x != 0: ret = str(x % base) + ret x //= base return ret | |
| Constant Value Error | Constant value error | ⬇ # [Task 51] SantaCoder def remove_vowels(text): vowels = ’aeiou’ return ’’.join([char for char in text if char not in vowels]) | ⬇ # [Task 51] Ground Truth def remove_vowels(text): return "".join(list(filter(lambda ch: ch not in "aeiouAEIOU", text))) |
| Reference Error | Wrong method/variable | ⬇ # [Task 101] GPT-3.5 def words_string(s): words = s.split(", ") words = [word.split(" ") for word in words] return words | ⬇ # [Task 101] Ground Truth def words_string(s): words = (s.replace(",", " ")).split() return [word for word in words if word != ""] |
| Undefined name | ⬇ # [Task 83] CodeGen-16B def starts_one_ends(n): if n == 1: return 1 return starts_one_ends(n-1)+ends_one_starts(n-1) | ⬇ # [Task 83] Ground Truth def starts_one_ends(n): if n == 1: return 1 return 18 * 10 ** (n - 2) | |
| Operation/ Calculation Error | Incorrect arithmetic operation | ⬇ # [Task 44] InCoder-1B def change_base(x: int, base: int): digits = [] while x: digits.append(str(x % base)) x /= base return ’’.join(reversed(digits)) | ⬇ # [Task 44] Ground Truth def change_base(x: int, base: int): if x == 0: return "0" ret = "" while x != 0: ret = str(x % base) + ret x //= base return ret |
| Incorrect comparison operation | |||
| Garbage Code | Only comments | ⬇ # [Task 152] InCoder-1B def compare(game,guess): # Regenerate the task description. | ⬇ # [Task 152] Ground Truth def compare(game,guess): return [abs(game[i] - guess[i]) for i in range(len(game))] |
| Meaningless code snippet | ⬇ # [Task 138] StarCoder def is_equal_to_sum_even(n): pass | ⬇ # [Task 138] Ground Truth def is_equal_to_sum_even(n): return n >= 8 and n % 2 == 0 | |
| Wrong (logical) direction | ⬇ # [Task 20] InCoder-1.3B def find_closest_elements(numbers): closest_to_one, closest_to_two = numbers[0], numbers[1] for number in numbers: if number < closest_to_one: closest_to_one = number if number > closest_to_two: cloeset_to_two = number return closest_to_one, closest_to_two | ⬇ # [Task 20] InCoder-1.3B def find_closest_elements(numbers): min_diff, min_pair = float("inf"), None for l, r in zip(numbers[:-1], numbers[1:]): diff = r - l if diff < min_diff: min_diff = diff min_pair = (l, r) return min_pair | |
| Incomplete Code/ Missing Steps | Missing one step | ⬇ # [Task 16] InCoder-1B def count_distinct_chars(string): return len(set(string)) | ⬇ # [Task 16] Ground Truth def count_distinct_chars(string): return len(set(string.lower())) |
| Missing multiple steps | |||
| Memory Error | Infinite loop | ⬇ # [Task 100] CodeGen-16b def make_a_pile(n): if n % 2 == 0: return [n] + make_a_pile(n+2) else: return [n] + make_a_pile(n+1) | ⬇ # [Task 100] Ground Truth def make_a_pile(n): ans, num = [], n for _ in range(n): ans.append(num) num += 2 return ans |
3.1.1. Semantic Characteristics
Condition Error includes missing condition and incorrect condition. Missing condition is when necessary conditions are omitted, while incorrect condition is when conditions are incorrectly formulated, leading to errors.
Constant Value Error is an error that occurs when an incorrect constant value is set, which can occur in function arguments, assignments, or other parts of the code.
Reference Error involves incorrect references to variables or functions, which includes the usage of an incorrect function or variable that does not match the requirement (wrong method/variable) and reference to a variable or method name that has not been defined (undefined name).
Operation/Calculation Error indicates the mistakes in mathematical or logical operations.
Garbage Code is defined as unnecessary or irrelevant code that does not contribute to the intended functionality. It can occur in several forms: a meaningless code snippet, where the code, though syntactically correct, is irrelevant to the assigned task; only comments, where the code consists exclusively of comments without any executable statements to address the task; or wrong (logical) direction, where the code significantly deviates from the intended task logic and expected outcomes.
Incomplete Code/Missing Steps indicates the absence of crucial steps to achieve the task.
Memory Error includes infinite loop, which is a loop or recursion that never terminates.
Comparison between LLMs. Fig. 1 shows the distribution of the 13 semantic characteristics for each LLM. On the one hand, we find that several characteristics are frequently shared among all LLMs, such as incorrect condition and wrong (logical) direction. This implies that all LLMs struggle with certain kinds of task requirements, such as handling complex logic conditions, regardless of their model size and capability.
On the other hand, small models such as InCoder and CodeGen are more likely to generate meaningless code and code that miss multiple steps, while larger models such as GPT-3.5 and GPT-4 tend to make more constant value errors and arithmetic operation errors. Notably, incorrect code generated by GPT-4 only exhibited 9 of the 13 semantic characteristics, while incorrect code generated by smaller models exhibited all sorts of errors. One plausible explanation is that GPT-3.5 and GPT-4 are much larger and are thus better at interpreting task descriptions. For instance, both CodeGen-16B and InCoder-1.3B missed the subtle requirement in Example LABEL:exp3 (Task 85) that “the elements must be even numbers,” resulting in missing steps within the generated code. However, GPT-3.5 and GPT-4 were able to generate the correct solution successfully.
CodeGen-16B
InCoder-1.3B
GPT-3.5
GPT-4
Loop ErrorFor error
While error
Return Error Incorrect return value
Method Call Error Incorrect function name
Incorrect function arguments
Incorrect method call target
Assignment Error Incorrect constant
Incorrect arithmetic
Incorrect variable name
Incorrect comparison
Import Error Import error
Code Block Error Incorrect code block
Missing code block
Finding 1: The most common semantic characteristics among six LLMs are wrong (logical) direction and incorrect condition. Compared with the other four LLMs, GPT-3.5 and GPT-4 do not often generate code with missing steps.
Furthermore, we found that even for the same failure task, failed code generated by different LLMs could have different semantic characteristics. Example LABEL:exp-failure-case shows an example of three LLMs that made different semantic characteristics. In this example, CodeGen-16B misses a few steps of handling the “unit digits” (missing multiple steps), while InCoder-1.3B completely ignored the task (wrong logical direction). GPT-3.5 only made a small mistake with parentheses (incorrect arithmetic operation). One takeaway from such observation is that ensemble of different LLMs might improve the code generation success rate.
CodeGen-16B
GPT-3.5
InCoder-1.3B
We also observed similar patterns in the LOC of correct solutions. Specifically, we looked into programming tasks whose correct solution has more than 12 lines of code. We found that 55.9% of these failed tasks are Garbage Code. FFor example, Task 105 requires three steps: 1) sorting an array of integers between 1 and 9 inclusive, 2) reversing the sorted result, and 3) replacing each digit with its corresponding name as a string. The correct solution has 24 lines of code. However, CodeGen-16B failed to understand the task requirements and returned only an empty array.
3.2. RQ4: How does code length relate to the types of LLMs’ code generation errors?
We did not find any statistical evidence that the generated code length is related to a specific task’s success. As shown in Fig. 2, in most cases, the passed tasks’ code length is similar to the failed tasks’ code length on all six LLMs. Mann-Whitney -test (Mann and Whitney, 1947) results further suggest that the mean differences are not statistically significant in all cases (). Finding 3: There is no statistical evidence indicating that there is a difference between the passed tasks’ code length and the failed tasks’ code length.
Though the average lengths of the code generated by CodeGen-16B, InCoder-1.3B, GPT-3.5, GPT-4, SantaCoder, and StarCoder (29.3, 32.3, 25.2, 29.6, 26.5, and 15.4 words, respectively) are short, we noticed a small portion of the generated code with a large length (i.e., longer than 100 words). Therefore, we further investigated these long code snippets generated by LLMs. In particular, we found that out of 36 code snippets with a length longer than 100 words, only 4 of them were correct. These results indicate that when the generated code is extremely long, there is a small chance that the code is correct. We further dug into those failure cases. Specifically, we found that 28 out of the 32 failed code snippets (87.5%) are Garbage Code, where 22 are Meaningless code snippet, 4 are Only comments, and only 2 are Wrong (logical) direction. Regarding Meaningless code snippet, we manually checked the 22 code snippets and found that all of them encountered the same problem: repeatedly generating similar lines/blocks of code. For instance, Task 6 requires the program to output the deepest level of nested parentheses. However, SantaCoder repeatedly generated the code to replace the parentheses with empty strings.
3.3. RQ4: Does partially failed code exhibit different characteristics compared with fully failed code?
Since HumanEval (Chen et al., 2021) provides test cases for each task, we are interested in whether completely failed code (i.e., failing all test cases) have different error characteristic patterns compared with partially failed code (i.e., failing a subset of test cases). Among the 557 incorrect solutions collected in Sec. 2.3, 243 are completely failed code and 314 are partially failed code. Fig. 3 shows the test pass rate distribution of incorrect solutions with different semantic characteristics. Only comments and meaningless code are the top two characteristics that lead to complete failures. Though undefined name sounds like a small error, the majority of code with undefined names also leads to complete test failures due to runtime crashes. Surprisingly, we found that while some error characteristics, such as missing multiple steps, sound severe by definition, they do not often lead to complete failures. After digging into some instances, we noticed that this is because LLMs do not completely misunderstand the task description, and the generated code can still pass some weak test cases. For instance, Task 125 requires the program to split on commas if there is no space in the string. If there is also no comma in the string, the program should perform other operations. However, the code generated by CodeGen-16B only split the input string into spaces while missing the remaining steps. As a result, it can only pass the test cases that include spaces, leading to a partially failed task.
Another surprising finding is that code with the wrong direction can pass some test cases accidentally. For example, in Task 75, the code generated by InCoder-1.3B does not follow the task instructions. However, it can pass a few test cases, such as 5, 10, and 30. One possible reason is that since the prompt includes test cases, LLMs may have memorized a superficial correlation between test cases and some other irrelevant solutions that pass those test cases.
In terms of syntactic characteristics (Fig. 4), we observed an obvious difference on if error, where the majority of code with this error led to partial failure (82% v.s. 18%). This is because when encountering if error, LLMs were more likely to only misinterpret a specific condition requirement while generating correct code for the other parts. For instance, in Task 0, SantaCoder only considered the adjunct elements in the given list numbers. As a result, it failed to pass the test case where the two closet elements were not adjunct.
4. Discussion and Future Work
The findings of our study shed light on several future directions to improve the quality and reliability of LLM-generated code.
Repairing Errors in LLM-generated Code. Our Finding 3.1.1, Finding LABEL:finding:2, and Finding LABEL:finding:3 reveal that errors made by different LLMs exhibit a wide range of different semantic and syntactic characteristics. Single-line logic errors such as if errors and incorrect function arguments may be relatively easy to fix, since existing automated program repair (APR) techniques are specialized to fix such errors (Xiong et al., 2017; Saha et al., 2017; Mechtaev et al., 2016; Ghanbari et al., 2019; Jiang et al., 2018; Wen et al., 2018; Liu et al., 2019; Li et al., 2022; Ye et al., 2022). However, our Finding LABEL:finding:3 indicates that many errors made by code LLMs involve multiple lines of code, which require substantial edits to fix. To overcome the limitations of traditional APR techniques, recent work has proposed to leverage LLMs to build program repair agents that are capable of synthesizing more complex patches (Fan et al., 2023; Zhang et al., 2024; Xia et al., 2024). These approaches typically rely on coarse-grained information such as error messages to guide the repair process. By contrast, our taxonomy provides a more detailed analysis of code generation errors based on their semantic and syntactic characteristics. We believe that leveraging this fine-grained information to repair errors in LLM-generated code is a promising direction. Specifically, one can first train a machine learning model to predict the types of errors. Then, the predicted errors can then be encoded in the prompt to guide the repair agent. For instance, instead of prompting the repair agent with “fix the bug at Line 6,” one can prompt the agent with “fix the incorrect function arguments at Line 6,” which may lead to a more accurate patch.
Fault Localization for LLM-generated Code. Precisely locating the error location is an important first step for fixing code generation errors. As shown by Finding LABEL:finding:3, code generation errors can occur in a variety of code constructs, which poses challenges to locating them precisely. Although many fault localization approaches have been proposed (Wen et al., 2019; Abreu et al., 2009; Li et al., 2019, 2021), these methods may still fall short in locating errors in LLM-generated code. Traditional approaches, such as spectrum-based fault localization, rely heavily on high-quality test cases that comprehensively examine different execution paths in a program. However, it is effortful and time-consuming to design test cases, which limits the utility of these approaches. Recently, deep learning has also been applied to fault localization (Li et al., 2019, 2021), typically using features extracted from source code and error messages. For code generated by LLMs, there is richer information from different sources, such as logits, token probability distribution, and self-attention scores at each decoding step. It would be interesting to investigate whether such information, together with error messages and other code characteristics, can be leveraged to predict at which step the model may start generating incorrect code.
Estimating the Correctness of LLM-generated Code. Our Finding 3.1.1 shows that current LLMs continue to struggle with understanding and solving complex task requirements. This observation raises several interesting research questions. First, for a given code generation LLM, can we develop a method to estimate the upper bound of task complexity that this LLM can handle? In this study, we have only experimented with simple metrics such as prompt length and LOC of the correct solution. Further investigation on this question could help assess the capabilities and limitations of a code generation LLM more accurately, leading to more reliable and trustworthy code generation. Second, for a given programming task, can we use prompt characteristics, such as length, to predict the task’s difficulty and the correctness of the generated code? This could potentially benefit software engineering tasks such as code review. For example, such estimations could help developers allocate more time to reviewing code that is likely to be less correct and associated with higher task difficulty.
5. Related Work
5.1. LLM-based Code Generation
In recent years, Large Language Models (LLMs) have not only demonstrated great performance in various natural language processing tasks (Dong et al., 2019; cha, 2023), but also shown great potential in code-related tasks, such as generation (Chen et al., 2021; Fried et al., 2022; Nijkamp et al., 2022; Wang et al., 2023; Zhu et al., 2024), summarization (Ahmad et al., 2020; Niu et al., 2022; Cai et al., 2024), understanding (Wang et al., 2022; Ding et al., 2024b), search (Feng et al., 2020; Zhang et al., 2022b; Jain et al., 2023) and translation (Roziere et al., 2020). As such, several LLMs specifically trained or fine-tuned on code data have been introduced, e.g., AlphaCode (Li and Choi, 2022), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023a), CodeGen (Nijkamp et al., 2022) and Qwen (Bai et al., 2023). A notable example is CodeLlama (Roziere et al., 2023), an adaptation of Llama-2 (Touvron et al., 2023). To enhance its code understanding and generation, CodeLlama introduced three significant modifications: (1) the collection of higher-quality code data for training, (2) the implementation of Fill-in-the-Middle (FIM) task (Fried et al., 2022) and (3) the adoption of instruction fine-tuning to better respond to code-related queries. Additionally, CodeLlama employed the self-instruct technique, where models generate programming questions and answers that are subsequently used for fine-tuning. Follow-up work can then be roughly categorized into three groups: (1) those that focus more on collecting more diverse, high-quality code-related training data, (2) those that seek better instruction fine-tuning techniques, and (3) those that utilize existing LLMs and devise better prompting strategies (Zhang et al., 2023; Bairi et al., 2024; Gao et al., 2023; Chen et al., 2023; Olausson et al., 2023; Li et al., 2023b; Ding et al., 2024a). For the first line of work, StarCoder 2 (Lozhkov et al., 2024) built the Stack v2 dataset, which is 4x larger than its predecessor, with up to 4.3 trillion tokens from more diverse sources such as GitHub pull requests, Kaggle notebooks, and code documentation. Additionally, DeepSeek-coder (Guo et al., 2024) introduced an approach for analyzing dependencies between code files using topological sorting, aggregating linked code lines into a single training sample, enabling LLMs to understand repository-level code bases better. For the second line of work, a representative example is WizardCoder (Luo et al., 2024), which implemented an “instruction evolution” strategy, where LLMs are iteratively guided to generate increasingly complex instructions covering a broader range of topics and skills. Later, MagicCoder (Wei et al., 2023) introduced OSS-Instruct, a dataset where GPT-3.5 Turbo is prompted to create coding problems and their solutions based on collected code snippets. OSS-Instruct consists of 75k fine-tuning instances, and the resulting MagicCoder, based on Code Llama, has demonstrated superior performance over GPT-3.5 on the HumanEval benchmark. The third line of the work focuses on designing more complex prompting strategies based on existing pret-trained LLMs. For example, Self-Debugging (Chen et al., 2024) enables LLMs to debug their own code by prompting with execution results and instructions for code explanations, demonstrating better accuracy than GPT-3.5 and GPT-4 on the MBPP dataset (Austin et al., 2021). Self-Refine (Madaan et al., 2024) iteratively enhances code generation by instructing the LLM to provide feedback on its outputs and refine them, achieving up to 13% improvement on CodeX (Chen et al., 2021). Cycle (Ding et al., 2024a) also leverages a similar idea of self-refinement based on the execution results of test suites. Different from Self-Debugging and Self-Refine, the Cycle includes a fine-tuning stage to enhance LLMs’ abilities to refine faulty programs. As a result, Cycle outperforms models with 3 times more parameters in self-refinement. CodeT adopts a test-driven prompting method (Chen et al., 2023). Rather than relying on existing test suites, CodeT instructs LLMs to simultaneously generate multiple code snippets and their corresponding tests, selecting the final output based on the highest consensus. SKCoder (Li et al., 2023b) utilizes a retriever-based pipeline to extract code sketches from existing databases. Distinct from these efforts that aim to train better LLMs, our research focuses on analyzing the errors LLMs produce. We aim to develop a taxonomy of these errors, with the hope that our findings will assist the development of new methods in the direction and contribute to better LLM-enabled intelligent software engineering.
5.2. Quality of LLM-generated Code
Evaluating the quality of LLM-generated code can reveal the current approach’s shortcomings and guide future improvement. While prior research has delved into various facets such as robustness (Arakelyan et al., 2023; Liu et al., 2024b), security (Pearce et al., 2022; Steenhoek et al., 2024), and usability (Barke et al., 2023; Al Madi, 2022), studies specifically examining the correctness of LLM-generated code (Chen et al., 2021; Nguyen and Nadi, 2022; Jesse et al., 2023; Liu et al., 2023, 2024a) are of particular relevance to our work. Liu et al. (Liu et al., 2023) evaluated the quality of code generated by ChatGPT, addressing various factors, including compilation and runtime errors, output correctness, coding style, maintainability, and performance. Their work, however, focused more on concrete error types. By contrast, our work focuses on code generation. We investigate a broader range of LLMs and examine the error characteristics’ correlations with prompt complexity, code similarity, etc. Pan et al. (Pan et al., 2024) introduced a taxonomy centered on code translation bugs. These bug types include code translate-specific problems (e.g., semantics differences between different programming languages) and logic and data-related bugs (e.g., the inclusion of logic not in source code and mismatched output data structure), which share some similarities with us. However, there is a key distinction in the origin of errors where Pan et al.’s work focuses on code-to-code translation, while our focus is on the issues in LLMs’ code generation from natural language specifications. Finally, Liu et al. (Liu et al., 2024a) conducted a study of the quality of code generated by ChatGPT, assessing their correctness, understandability, and security. In their examination of code correctness, they primarily focused on compile errors and runtime crashes. Our research differs from theirs in two key respects: (1) Our study subjects are more diverse with both open-source and closed-source models; (2) our taxonomy also considers behavior deviations informed by test failures in addition to compiling errors and crashes. Based on the findings, we further provide actionable suggestions and implications for future enhancements in LLM-enabled code generation.
5.3. Taxonomy on Software Defects
In software engineering, collecting, classifying, and analyzing software defects has been recognized as an important research topic for improving software quality. Building a systematic taxonomy of software bugs can help stakeholders understand the pitfalls of target systems and guide developers for better development practices. One of the early attempts along the direction was the orthogonal defect classification (ODC) proposed by IBM Research (Chillarege et al., 1991, 1992). Since then, numerous endeavors have been made to construct defect taxonomies from various angles. These include investigations targeting different programming languages, such as C (Le Goues et al., 2015; Tan et al., 2017), Javascript (Hanam et al., 2016; Gyimesi et al., 2019), Java (Pan et al., 2009), Python (Zhang et al., 2022a), Julia (Rahman et al., 2023), and multi-language programs (Lin et al., 2017; Kochhar et al., 2016). Others delve into different applications, such as Infrastructure as Code (IaC) (Rahman et al., 2020) and deep learning stacks (Humbatova et al., 2020). Different from these attempts, our work mainly focuses on building a systematic taxonomy for LLM-generated code.
6. Threats to Validity
Threats to Internal Validity. One potential threat comes from our manual analysis process, where labelers may have different opinions and sometimes may make mistakes. To mitigate this, four of the authors first performed open coding and iteratively refined our codebook until a substantial agreement was achieved before two of the authors labeled the complete dataset. Our final Fleiss’ Kappa regarding the semantic and syntactic characteristics are 0.91 and 0.91, respectively, indicating almost perfect agreement (Landis and Koch, 1977).
Threats to External Validity. One potential threat lies in the choice of dataset. Considering the extensive labeling effort (e.g., running the code in an IDE and performing step-by-step debugging to locate the root cause), we have only labeled one dataset. In future work, one may consider labeling more datasets to confirm the generalizability of our findings. Nevertheless, given the size of our labeled dataset (557 code snippets), we believe our findings can be generalized to other similar datasets, such as MBPP (Austin et al., 2021) and APPS (Hendrycks et al., 2021). Moreover, our study only covers Python programs, which might not generalize to programming languages that are very different from Python (e.g., PHP and Rust). We chose Python because it is one of the most popular object-oriented programming languages. In future work, we plan to expand our study with programming tasks from other languages. Finally, we have only experimented with six LLMs in this study due to the intensive labeling effort (i.e., 328 person-hours for labeling 557 incorrect code snippets). Given that code generation with LLMs is a fast-developing research field, we also plan to expand our study by labeling errors from more recent code LLMs, e.g., MagiCoder (Wei et al., 2023) and CodeLlama (Roziere et al., 2023).
Threats to Construct Validity. In RQ3, we use the length of the task description and the length of the correct solution to estimate the task complexity. Although they may not be the best metrics for such estimation, we believe this is a pragmatic choice since there is no commonly used metric for measuring task complexity for LLM’s code generation. In future work, researchers may consider designing new metrics to estimate the task complexity and investigate its correlation with LLM’s code generation capabilities.
7. Conclusion
In this paper, we present an empirical study on code generation errors made by large language models. We first derived a taxonomy of LLMs’ code generation errors based on six popular LLMs’ failure cases on the HumanEval dataset (Chen et al., 2021) through open coding and thematic analysis. We labeled a total of 557 errors committed by these code LLMs according to the established taxonomy. We found that these LLMs exhibited different distributions of semantic and syntactic error characteristics. We further analyzed the bug-fixing effort, the impact of task complexity, and the correlation between test pass rates and different kinds of errors. At the end of the paper, we discuss the implications of our study and propose future research opportunities for improving the quality and reliability of code LLMs.